PROGRAMS FOR TWO-ROTOR PROBLEMS

\underline{PETER GRONER}, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110.

The simulation and fitting of spectra of molecules with two periodic large-amplitude motions (”internal rotations”) has been a challenging problem for some time. Three Fortran programs (for PC) treat different aspects of the rotation-internal motion Hamiltonian. TACIR (Two Asymmetric Coupled Internal Rotors) solves the internal motion (J = 0) Hamiltonian and predicts torsional transitions for equivalent or nonequivalent internal rotors of any periodicity from internal rotation ”constants” and potential, each defined as two-dimensional Fourier series. It is a significantly expanded version of a program described earlier ^{a}. The coefficients of the potential and other spectroscopic constants can be adjusted by the least-squares method to fit energy differences (transition frequencies and tunneling splittings). It has been applied recently to 3-methyl-1,2-butadiene, (CH_{3})_{2}C=C=CH_{2} ^{b}, and acetone, (CH_{3})_{2}C=O ^{c}. Program ERHAM (Effective Rotational HAMiltonian) solves the effective rotation-internal motion Hamiltonian for molecules with equivalent internal motions d. It predicts rotational transitions and intensities and fits spectroscopic constants (by least-squares) simultaneously for several but uncoupled (bound) vibrational states ^{e}. The third program, TWOSER (TWO-dimensional fourier SERies) determines the 2-D Fourier series of the internal rotation ”constants” used in TACIR for asymmetric internal rotors from the structural parameters of several conformations.

aP. Groner and J. R. Durig, J. Chem. Phys. 66, 1856 (1977).

bS. Bell, P. Groner, G. A. Guirgis and J. R. Durig, J. Phys. Chem. A, 104, 514 (2000).

cP. Groner, J. Mol. Struct. 550-551, 473 (2000).

dP. Groner, J. Chem. Phys. 107, 4483 (1997).

eP. Groner, S. Albert, E. Herbst and F. C. De Lucia, Astrophys. J. 500, 1059 (1998).